OpenStudy (honeymonster):

can you please check my answer for finding x and y in this simultaneous equation..posted in replies

OpenStudy (honeymonster):

\[x+2y=3\] \[x+y =2\] 0x+y=1 x+y=2 2-1=1=x y=1 x=1

OpenStudy (anonymous):

yes that's right

OpenStudy (honeymonster):

thank you...